ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 51, 54, 5, 52 Zoch, Avery, S., University of Houston-Downtow, U. S. A. The Jordan Polygonal Theorem In this article it is proven that the jordan curve theorem can be extended to the case of a polygon with positive area. There is an interesting use of the Erdos-Szekeres theorem along the way to this result and a corollary that has proved very useful in other works of the author that involves the Euler characteristic theorem for planer simple graphs. Reference: A. S. Zoch, The Jordan Polygonal Theorem, American Mathematical Society (current review)